# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v2

@@ 5.3 Potential improvements and future work
The reliability of navigation and the smoothness of the trajectory could be further improved by investigating alternative camera systems and image processing methods (e.g., image super-resolution) with a better capability of detecting thin branches. Blindsides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or storing historical data to occupancy grid maps in addition to the currently visible area.
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the takeoff location of the drone, limiting the low-altitude flying to relatively flat areas. To comprehensively map uneven forest plots, the altitude limits should be modified to be dynamic and concerning the local terrain level.
Drift compensation remains an important objective for future development. Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. The potential methods for GNSS-Fusion include using the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g. GVINS (Cao, Lu, and Shen 2022). Even though the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forests.
@@
In this study, the performed flight paths were straight flights through the sample plots, and the DBH estimation was performed only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides. Reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the higher DBH estimation error obtained with the trees with large diameters. The global path covering the whole test plot could be implemented either with a simple waypoint generation algorithm as Liang et al. (2024) or with more advanced exploration methods such as proposed by Zhou et al. (2021a) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage.
While this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications. Including an additional lightweight camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds. With the developed camera-based prototype, data collection extends beyond point clouds to encompass applications such as image-based mapping and object detection. This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damages caused by insect pests.
The prototype also opens possibilities for research of real-time forest characteristics analysis. In this study, the point clouds were obtained through post-processing. However, the onboard Intel Realsense D435 camera is also capable of generating 3D point clouds in real time. Additionally, image-based object detection can be performed in real-time, e.g., with YOLO (Redmon et al. 2016). While only computationally light analysis can be conducted on the onboard computer, real-time analysis with higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.

### v3

@@ 5.3 Potential improvements and future work
The reliability of navigation and the smoothness of the trajectory could be further improved by investigating alternative camera systems and image processing methods (e.g., image super-resolution) with a better capability of detecting thin branches. Blind sides in the obstacle avoidance map could be tackled either by sensors with 360° FOV (e.g., 360° camera or multi-camera system) or by storing historical data in the occupancy grid map in addition to the current visible area.
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas. To capture images at the DBH estimation altitude in uneven forest plots, the flight altitude limits should be modified to be dynamic with respect to the local terrain level.
The drift compensation remains also an important objective for future development. Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao, Lu, and Shen 2022). Although the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forest.
@@
In this study, the flight paths were straight flights through the sample plots, and the DBH values were estimated only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides. In the study by He et al. (2025), the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed. This suggests that reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the DBH estimate errors obtained with the system utilized in this study. In particular, this is expected to improve the quality of the estimation for trees of large diameter that yielded a higher RMSE in this study than trees with a DBH of less than 30 cm. The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as Liang et al. (2024), or with more advanced exploration methods such as those proposed by Zhou et al. (2021) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage. In potential follow-up studies, the DBH estimation potential of the system could also be evaluated in alternative forest environments, for example, in sparser forest environments.
Although this study focused on demonstrating the potential for forest characteristics estimation through DBH estimation and tree detection, the scope of data collection extends well beyond these applications. Including an additional lightweight camera pointing upwards would enable a more comprehensive reconstruction of the tree trunks to perform also stem curve estimation based on the point clouds. With the developed camera-based prototype, data collection extends beyond point clouds to encompass applications such as image-based mapping and object detection. This opens up a variety of potential uses, including classifying and counting tree species and other vegetation, as well as monitoring bark damage caused by insect pests.
The prototype also opens up possibilities for the study of real-time forest characteristics analysis. In this study, the point clouds were obtained through post-processing. However, Intel RealSense D435 is also capable of generating 3D point clouds in real time. Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon et al. 2016). While only computationally light analysis can be conducted on the onboard computer, real-time analysis with a higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.